OpenStudy (anonymous):

Can someone help and explain this equation for me? 1/7 (14 - 7p) - 2 = -2 (½ p + 3) + 6

OpenStudy (prizzyjade):

\[\frac{ 1 }{ 7 }(14-7p)-2 = -2 (\frac{ 1 }{ 2 }p +3 ) +6\]

OpenStudy (prizzyjade):

\[(2-p)-2=(-p-6)+6\] 2-p-2=-p-6+6 -p=-p -p+p=0 0=0